Summary and Descriptive Statistics

There is often the requirement to evaluate descriptive statistics for data within the organization or for health care information. Every year the National Cancer Institute collects and publishes data based on patient demographics. Understanding differences between the groups based upon the collected data often informs health care professionals towards research, treatment options, or patient education.

Using the data on the "National Cancer Institute Data" Excel spreadsheet, calculate the descriptive statistics indicated below for each of the Race/Ethnicity groups. Refer to your textbook and the Topic Materials, as needed, for assistance in with creating Excel formulas.

Provide the following descriptive statistics:

  1. Measures of Central Tendency: Mean, Median, and Mode

  2. Measures of Variation: Variance, Standard Deviation, and Range (a formula is not needed for Range).

  3. Once the data is calculated, provide a 150-250 word analysis of the descriptive statistics on the spreadsheet. This should include differences and health outcomes between groups.

APA style is not required, but solid academic writing is expected.
Answer #1

Answer:

Fomulas for computations are

Mean \(=\frac{\sum X i}{N}\)

Variance \(=\frac{\sum(X i-\bar{X})^{2}}{N}\)

Here \(\mathrm{N}=16\) which is even

Median \(=\frac{(n / 2) o b s+(n / 2+1) o b s}{2}\)

Rnage=Maximum-Minimum

American/Alaska Native


Mean43.275
Standard error1.31622883
Median43.85
Mode#N/A
Standard Deviation5.264915321
Sample Variance27.71933333
Kurtosis-0.136660029
Skewness-0.383344382
Range19.7
Minimum32
Maximum51.7
Sum692.4
Count16

Interpretation: From descriptive statistics, we say that in America/Native Alaska region Mean rate for Lung and bronchus cancer over the last 16 years is 43.275 per 100 with a standard deviation of 5.26 per 100, Variance of 27.71 per 100 with the maximum number of cases approx 52

Asian/Pacific Islander


Mean38.5125
Standard Error0.595810023
Median38.9
Mode36.6
Standard Deviation2.383240091
Sample Variance5.679833333
Kurtosis-0.712232536
Skewness-0.555307749
Range7.8
Minimum34
Maximum41.8
Sum616.2
Count16

Interpretation: From descriptive statistics, we say that in Asian/Pacific Islander region Mean rate for Lung and bronchus cancer over the last 16 years is 38.51 per 100 with a standard deviation of 2.38 per 100, Variance of 5.67 per 100 with a maximum number of cases approx 42.

Black (Includes Hispanic)


Mean70.06875
Standard Error1.685025
Median71.4
Mode#N/A
Standard Deviation6.740101
Sample Variance45.42896
Kurtosis-0.94965
Skewness-0.50796
Range21.6
Minimum57.4
Maximum79
Sum1121.1
Count16

Interpretation: From descriptive statistics, we say that in the white (includes Hispanic) region Mean rate for Lung and bronchus cancer over the last 16 years is 70.06 per 100 with a standard deviation of 6.74 per 100, Variance of 45.42 per 100 with a maximum number of cases approx 79.

Hispanic (Any Race)


Mean31.49375
Standard Error0.724538287
Median32.1
Mode34.1
Standard Deviation2.898153148
Sample Variance8.399291667
Kurtosis-0.88797923
Skewness-0.611418136
Range9
Minimum26
Maximum35
Sum503.9
Count16

Interpretation: From descriptive statistics, we say that in the Hispanic(Any Race)region Mean rate for Lung and bronchus cancer over the last 16 years is 31.49 per 100 with a standard deviation of 2.89 per 100, Variance of 8.39 per 100 with a maximum number of cases approx 26.

White (Inclusions Hispanic)


Mean62.725
Standard Error1.278720063
Median64.55
Mode65.8
Standard Error5.114880253
Sample Variance26.162
Kurtosis-1.087933179
Skewness-0.55685628
Range15.6
Minimum53.2
Maximum68.8
Sum1003.6
Count16

Interpretation: From descriptive statistics, we say that in the white (includes Hispanic) region Mean rate for Lung and bronchus cancer over the last 16 years is 62.725 per 100 with a standard deviation of 5.11 per 100, Variance of 26.16 per 100 with a maximum number of cases approx 68.8.

Conclusion: Form the descriptive statistics We can say that the region Black(Includes Hispanic) has the highest number of cases over the last 16 years with the greatest mean rates of cases among all regions and white (includes Hispanic) comes on the second number then American/Native Alaska, Asian/Pacific Islander and Hispanic (Any Race) in decreasing order has the cases of Lung cancer.


answered by: HSJ